OpenStudy (anonymous):

log10(18)-log10(3x)=log10(2) Someone help please?

OpenStudy (anonymous):

x=3

OpenStudy (anonymous):

can you tell me how you got that please?

OpenStudy (anonymous):

log10(18)-log10(3x) simplifies to log10(18/3x) sooo.. log10(18/3x)=log10(2) The bases are the same so drom the logs. 18/3x=2 18=2*3x 18/2=3x 9=3x 3=x

OpenStudy (anonymous):

can you help me with some more please?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

1/3 log 2(27)+log2(36)=log2(n)

OpenStudy (anonymous):

n=108

OpenStudy (anonymous):

can you explain it again? sorry ineed to understand this for my test

OpenStudy (anonymous):

http://www.rapidtables.com/math/algebra/logarithm/Logarithm_Rules.htm here's the list of logarithmic rules that you'll need to know then. Good luck on your test!

OpenStudy (anonymous):

thank you so much.!